# Cracking Vector Search Indexes

### v1

@@ 5.5. Control Mechanisms Ablation Study
(a) WHERE=OFF: QPS vs Recall
(b) WHERE=ON: QPS vs Recall
Figure 5. Comparison of QPS-Recall, showing effect of removing the ”where” to CRACK and REFINE mechanism
(a) Cumulative Time (OFF)
(b) Total Time Breakdown (OFF)
(c) Cumulative Time (ON)
(d) Total Time Breakdown (ON)
Figure 6. Comparison of cumulative time and total time breakdown, showing effect of removing the ”when” to CRACK and REFINE mechanism
CrackIVF uses two control mechanisms for the CRACK and REFINE build operations, as detailed in Section 4.2 and Section 4.3. The first mechanism, controlling where they should be applied, uses a set of heuristic rules. The second mechanism, controlling when they are executed, is a cost-budget-based approach combined with a predictive model to balance search and build times. This section demonstrates the importance of these controls in achieving optimal performance and additionally highlights the effect of changing the parameter default value $min\_pts=2$, which changes the heuristic rule “Don’t Crack: Rule 1“ (Section 4.2).
@@
Turning off the control mechanism for “where“: To illustrate the importance and accuracy of the heuristic rules for where to CRACK and REFINE, we switch them OFF and see the effect (Figure 5). Specifically, when “WHERE=OFF”, and no heuristic rule is used to decide where a crack or refine should happen, the index defaults to trying to apply these operations whenever there is enough budget for them (the “when“ mechanism is still ON). This seemingly random addition of new cracks and local refines leads to a final performance that is no better than the initial state of CrackIVF, before any query has been received (Figure 5(a)). On the other hand, when “WHERE=ON“, and our heuristic rules are applied to make decisions, CrackIVF converges to the Pareto optimal QPS-Recall trade-off performance (Figure 5(b)). In both examples, we send an equal number of queries to the index and in the same order.
Turning off the control mechanism for “when“:
@@
The budgeting mechanism, controlling when build operations are executed, helps balance the time spent on build vs search operations. By default, we use the parameter $a=0.5$, i.e., at most 50% of the total time may be spent on $CRACK$ or $REFINE$. The effect of turning this mechanism “WHEN=OFF“ is shown in Figure 6(a) and Figure 6(b). CrackIVF with the mechanism off defaults to cracking and refining after almost every query, and it only stops when the maximum number of partitions is reached (16,000). So even though the cracks and refines happen in regions that require them, they happen disproportionally often compared to the search overhead, leading to a final time spent on build operations that completely dominates the total time. On the other hand, when the mechanism is enabled (“WHEN=ON“), CrackIVF is able to efficiently balance the search and build costs (Figure 6(b) and Figure 6(c)). These runs are for the same 100k query trace.
Varying heuristic rule parameters:
(a) QPS vs Recall ($min\_pts=2$)
(b) QPS vs Recall ($min\_pts=32$)
Figure 7. Comparison of QPS vs Recall for different $min\_pts$ values.
CrackIVF’s heuristic rules use default parameters (Section 4.2). To avoid parameter tuning, we did not vary these values across experiments. However, testing revealed that the $min\_pts$ parameter can impact performance, dependent on the dataset. $min\_pts$ defines the minimum number of points that must be stolen by a new crack for it to be considered a “good“ and subsequently buffered. On the GloVe-25 dataset, changing $min\_pts$ from 2 to 32 notably improved the QPS-Recall performance of the final index (Figure 7). Originally, we set $min\_pts=2$, a very small value, to ensure we avoid adding empty cracks to the index. We assumed that any imbalances could be fixed by future $REFINE$ operations that redistribute points locally, which works for all datasets except for GloVe-25.
Although we can not make a definite statement, we hypothesize that this can be attributed to the fact that this dataset only has 25-dimensional embeddings. This leads to search times that are very low, and subsequently, our build operations budget grows slower. Thus, there are too many imbalances introduced with $min\_pts=2$, and a smaller budget for future $REFINE$ operations to correct them. By increasing $min\_pts$ to 32, we reduced the total number of local imbalances by avoiding to ever commit excessively small cracks, which in turn reduces the number of $REFINE$ needed to reach a good final performance.
A second hypothesis comes from closer inspection of the Figures, in which the better performing index converges to 10056 partition vs the default 16000. (Figure 7(a), Figure 7(b) ). But, we believe because GloVe-25 queries access the entire vector space evenly (Figure 2(b)) and it is a dataset with (1.18M points), we do not think a larger number of partitions hinders performance.
Nevertheless, we show that parameter values do matter across datasets, so index tuning can become a consideration. But for this work, across all the datasets we tested, outside of GloVe-25, the parameter defaults we have chosen have reliable performance.
Takeaways: Our results show that CrackIVF minimizes the cumulative costs across a large variety of individual datasets tested, but to fully grasp the potential performance benefits in embedding lakes, it’s essential to consider the aggregate across all datasets’ cumulative time savings and as the number of queries varies. For systems managing potentially thousands or millions of datasets, all saving are scaled equivalently, and choosing a separate index for each dataset is near-impossible, so the benefit from using CrackIVF is greater, as there is no one-size-fits all solution from the pre-built indexes.
@@
From our experiments, we make the following suggestion: Brute Force should be used as the default method up until a dataset surpasses 10–100 queries, at which point CrackIVF becomes the optimal choice. Even though CrackIVF can start with one partition to simulate a Brute Force approach, it has a slight memory and runtime overhead compared to it, from the setup time and storage of internal object structures and keeping track of the first few partition representatives. At scale, for embedding lakes with millions of datasets, even this small overhead can accumulate. Therefore, we recommend using CrackIVF once a dataset exceeds 100 queries and is thus likely to receive even more.
@@ 6. Related Work
Vector Databases: Vector Data Management Systems (VDBMS) (Wang et al., 2021; Weaviate, 2025; Pinecone, 2025; Zhang et al., 2023; Chen et al., 2024; Guo et al., 2022), are systems that support efficient ANN search, and have gained significant popularity in both research and industry. This coincides with the increase in popularity of Large Language Models (LLMs) (Brown et al., 2020), that use Retrieval-Augmented Generation (RAG) techniques (Lewis et al., 2020; Asai et al., 2023), to extend their capabilities beyond what they have been trained on, in order to answer queries based on external data sources. Vector Database Management System applications extend beyond the above use-case, as they are also used as platforms for semantic search (Salton et al., 1975; Mitra et al., 2018), recommendation systems (Li et al., 2017) and product search in e-commerce (Van Gysel et al., 2016). Given the importance of these operations, many existing data management systems are incorporating vector search capabilities (Corporation, 2025; pgvector contributors, 2025). Many of these systems also support extensions to dense vector ANN search, through specialized methods, allowing for specific query types such as filtered search (Gollapudi et al., 2023), sparse vector search (Formal et al., 2021), and hybrid search (Vespa.ai, 2024), which combines vector search with more traditional information retrieval techniques like BM-25 (Robertson et al., 2009). A recent survey of vector database management systems is found in (Pan et al., 2024).
@@
Vector Search Indexes: Approximate Nearest Neighbor (ANN) search is a well-studied field with decades of research, resulting in a large number of indexes. Broadly, the two dominant categories of ANN vector search indexes are partition-based and graph-based indices. Inverted index (IVF) based methods, which fall under partition based category, include SCANN and SOAR (Guo et al., 2020; Sun et al., 2024), which identify key mathematical properties of ANNS and use them to improve IVF index efficiency, SPANN (Chen et al., 2021), which shows how to leverage disk-based storage for search over billion-scale datasets along with supporting updates (Xu et al., 2023), and FAISS-IVF (Douze et al., 2024), an in-memory implementation utilizing GPUs (Johnson et al., 2019) and quantization techniques like PQ (Sivic and Zisserman, 2003; Jegou et al., 2010; Ge et al., 2013). In this work, we focus on IVF methods, and our cracking based approach could be extended for the above implementations. IVF indexes offer the best trade-off between index build time, search performance, and memory overhead, all of which are critical to balance for potential large scale deployments in embedding data lakes. Hash based methods (Andoni et al., 2015; Indyk and Motwani, 1998; Datar et al., 2004; Jain et al., 2008; Wang et al., 2017) and tree based methods (Muja and Lowe, 2014; Liu et al., 2004; Wang et al., 2013) also broadly fall under the partition-based index category. Finally, graph-based methods (Malkov and Yashunin, 2018; Malkov et al., 2014; Wang et al., 2012; Dong et al., 2011; Wang et al., 2013; Hajebi et al., 2011) currently deliver the state-of-the-art performance in terms of the Queries Per Second vs. Recall trade-off, but this comes at high memory and index construction cost. Thus, graph based indexes are for now ill-suited for scenarios where minimizing the index construction cost and low memory overhead is a important, for example in embedding lakes where a very large number of indexes is expected to be constructed.
@@
Adaptive indexes for specialized data types: Adaptive indexing techniques, such as Dumpy (Wang et al., 2023, 2024) and ADS (Zoumpatianos et al., 2014), have demonstrated significant effectiveness in data series similarity search. These indexes are specifically designed for time-series data and include methods like Symbolic Aggregate Approximation (SAX) (Lin et al., 2003) and indexable SAX (iSAX) (Shieh and Keogh, 2008). In our work, we focus on ANN indexing methods that are the standard used by RAG systems, and due to their broad applicability across various data modalities which lack sequential relationships. That said, specialized, adaptive and incrementally constructed indexes are relevant candidates for embedding lake deployments, particularly for the subset of the datasets with inherent sequential dependencies. This includes energy, weather and financial data, as well as audio and video recordings (Raza et al., 2015; Mirylenka et al., 2016; Shasha, 1999; Kashino et al., 1999).
@@
ANN index maintenance: Work on index maintenance focuses on mitigating performance degradation caused by data distribution shifts due to data updates (insertions and deletions). DeDrift (Baranchuk et al., 2023) handles data or query distribution shifts and identifies small and large partitions to re-cluster together, but it keeps the total number of partitions fixed. LIRE (Xu et al., 2023) performs localized split and merge operations when partitions grow too large or become too small after insertions and deletions. Ada-IVF (Mohoney et al., 2023) monitors partition size changes from updates and the partition access distribution from the queries. It combines this information to prioritize the maintenance operations that will improve index efficiency. Similar to us, these approaches demonstrate the effectiveness of localized maintenance operations for IVF indexes and the advantages of adapting to query distribution. However, they cannot be applied directly for our usage, which is incremental index construction as the number of queries increases. These works all trigger maintenance operations directly or implicitly from incoming updates (insertions/deletions). Thus they will never trigger in our static setting, whereas we do since we follow the database cracking paradigm, that states “index maintenance should be a byproduct of query processing, not of updates” (Idreos et al., 2007), and CrackIVF operations are triggered solely from the queries.
7. Conclusion and Future Work
@@
We introduced CrackIVF, a cracking-based IVF index for Approximate Nearest Neighbor Search (ANNS), aimed to be used with RAG systems over embedding data lakes. By dynamically adapting to increasing query workloads, it eliminates the need for costly up-front indexing, allowing immediate query processing. Our experimental results show that CrackIVF can process over 1 million queries before conventional methods even finish building an index, achieving 10-1000x faster initialization times. This makes it particularly effective for cold data, infrequently accessed datasets, and rapid access to unseen data. As future work, we want to adapt CrackIVF to handle query and data distribution shifts and identify reliable methods to dynamically set parameters across diverse datasets.
8. Acknowledgments

### v2

@@ 5.5. Control Mechanisms Ablation Study
This section is an ablation study exploring the control mechanisms “where” and “when” to apply $CRACK$ and $REFINE$.
@@
Turning off the control mechanism for “where”: To illustrate the importance and accuracy of the heuristic rules for where to CRACK and REFINE, we switch them OFF and see the effect (Figure 6). Specifically, when “WHERE=OFF”, and no heuristic rule is used to decide where a crack or refine should happen, the index defaults to trying to apply these operations whenever there is enough budget for them, since the “when” mechanism is still ON. This seemingly random addition of new cracks leads to a final performance that is only slightly better than the initial state of CrackIVF, before any query has been received (Figure 6(a)). On the other hand, when “WHERE=ON”, and our heuristic rules are applied to make decisions, CrackIVF converges to the Pareto optimal QPS-Recall trade-off performance (Figure 6(b)). In both examples, we test an equal number of queries and in the same order.
@@
Turning off the control mechanism for “when”: The budgeting mechanism, controlling when build operations are executed, helps balance the time spent on build vs search operations. By default, we use the parameter $\alpha=0.5$, i.e., at most 50% of the total time may be spent on $CRACK$ or $REFINE$. The effect of turning this mechanism “WHEN=OFF” is shown in Figure 7(a) and Figure 7(b). CrackIVF with the mechanism off defaults to cracking and refining after almost every query, and it only stops when the maximum number of partitions is reached (16,000). So even though the cracks and refines happen in regions that require them, they happen disproportionally often compared to the search overhead, leading to a final time spent on build operations that completely dominates the total time. On the other hand, when the mechanism is enabled, “WHEN=ON”, CrackIVF is able to efficiently balance the search and build costs (Figure 7(c) and Figure 7(d)). These runs are for the same 100k query trace. The parameter controlling the “when” mechanism is $\alpha$. The effect of varying $\alpha$ is shown in Figures 8(a) and  8(b). If build is 25% of the total time then it is too restrictive. On the other hand, if 75% of the total time can go to build operations, a very large number of refines keep triggering. Figure 8(b) illustrates the importance of monitoring the new buffered cracks rate and the refine regions in order to converge the index. After $\approx$50k queries, no new cracks occur, preventing further index growth. Refines also repeatedly trigger in the same regions, as shown by the striated pattern, offering no additional benefit, since the median response time does not improve after the refines. This indicates the index has converged. We monitor when this state is achieved and then converge the index, avoiding further build operations.
(a) SIFT1M $\alpha=0.25$
(b) SIFT1M $\alpha=0.75$
(c) Deep10M min_pts=32min\_pts=32
(d) SIFT1M min_pts=32min\_pts=32
(e) lastfm pts_crack_thr=16pts\_crack\_thr=16
(f) Glove50 pts_crack_thr=128pts\_crack\_thr=128
(g) Deep10M cv_max=8cv\_max=8
(h) SIFT10M cv_max=8cv\_max=8
(i) SIFT1M size_prctlsize\_prctl OFF
(j) Glove100 size_prctlsize\_prctl OFF
Figure 8. Ablation study of parameters.
Table 4. Parameters defining the Decision Rule Boundaries.
Parameter Restrict “good candidates” Tested Values
min_ptsmin\_pts 1, 2, 32
pts_crack_thrpts\_crack\_thr 16, 64, 128
cv_maxcv\_max 1, 2, 8
size_prctlsize\_prctl 1, 10, 25
Ablation on decision boundary parameters: Each parameter defines a decision boundary (Section 4.2). Changing it in one direction, makes the model more restrictive, classifying fewer candidates as good, while in the opposite direction it is more permissive. Table 4 indicates the restrictive direction and tested values (default in bold). We did multiple runs for each dataset and parameter combination. For each parameter we pick the most interesting results, showcased in Figure 8. We only vary one parameter at a time, the others are fixed to their default value. We found the model to be robust to changing a single parameter. Out of the 48 edge-case parameter-dataset combinations, 15 had a noticeable effect on the final index. We attribute the robustness to each heuristic involving two decision rules (Section 4.2), thus two parameters. When one is changed, the other still provides a reasonable boundary. Setting both parameters to restrictive extremes prevents any cracking, while setting both to relaxed extremes is equivalent to WHERE=OFF (Figure 6(a)), where every query is classified as good and the operations occur when the cost budget allows.
For min_ptsmin\_pts, we have the most consistent effect across all our observations. When we set it to a very restrictive value ($32$), the number of good cracks drops significantly, and thus the index converges to a smaller size, leading to lower performance (Figure 8(d)). This negatively affected all 6 datasets.
The pts_crack_thrpts\_crack\_thr, mostly affects the highest skew datasets. For example, on Last.fm performance worsens when the parameter is set to a permissive value, as this leads to over-partitioning of the region, as more cracks are allowed in a region (Figure 8(e)). But, the above value did not affect any other dataset. On the other hand, restricting it negatively affected 2 out of 6 datasets (Figure 8(f)). This was the same effect as that of min_ptsmin\_pts. When the rule is too restrictive, not enough cracks pass it so index can not fully grow.
Restricting cv_maxcv\_max can have a positive effect, e.g. on SIFT10M (Figure 8(h)) and on DEEP10M (Figure 8(g)). This leads us to believe that in some datasets allowing a higher local imbalance, could have beneficial overfitting effects to the query distribution. The optimal value for cv_maxcv\_max is not consistent across datasets, two other datasets had worse results at a high value, and our default values provide the most robust results.
All the values tested for size_prctlsize\_prctl had a negligible effect. Thus, to show the benefit of including the decision rule which size_prctlsize\_prctl controls, we switch it off entirely in Figures 8(i) and 8(j). Disabling it greatly affects final index. While the index grows, and it does not converge to the optimal state, as fewer refines occur.
Overall, we clearly see dataset-dependent effects which can change the optimal value of each parameter per dataset. This leaves room for further tuning and “learned” decision boundaries instead of setting them manually. Tuning of parameters exists in any indexing technique, the more you know about your dataset the better you can tune the index for it. We did not pursue tuning in this work, as the default values empirically provide very robust performance.
@@
From our experiments, we make the following suggestion: BruteForce should be used as the default method up until a dataset surpasses 10–100 queries, at which point CrackIVF becomes the optimal choice. Even though CrackIVF can start with minimal initialization cost, it is more complex and has a slight memory and runtime overhead compared to BruteForce due to the setup time and storage of internal object structures and keeping track of the first few partition representatives. At scale, in embedding lakes even a small overhead can accumulate. We recommend using CrackIVF once a dataset exceeds 100 queries and is likely to receive more.
@@ 6. Related Work
Vector Databases: Vector Data Management Systems (VDBMS) (Wang et al., 2021; Weaviate, 2025; Pinecone, 2025; Zhang et al., 2023; Chen et al., 2024; Guo et al., 2022; Deng et al., 2025; Kuffo et al., 2025; Xu et al., 2025; Cai et al., 2024) support efficient ANN search, and have gained significant popularity. This coincides with the increase in use of Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Asai et al., 2023), to extend the capabilities of LLMs beyond what they have been trained on. VDBMS applications also include semantic search (Salton et al., 1975; Mitra et al., 2018), recommendation systems (Li et al., 2017) and product search (Van Gysel et al., 2016). Many relational engines are incorporating vector search (Corporation, 2025; pgvector contributors, 2025). Many of these systems support extensions to dense vector ANN search, allowing for specific query types such as filtered search (Gollapudi et al., 2023), sparse vector search (Formal et al., 2021), and hybrid search (Vespa.ai, 2024). A recent survey covering vector database management systems can be found in (Pan et al., 2024).
@@
Vector Search Indexes: Approximate Nearest Neighbor (ANN) search has been well-studied. The two dominant categories of ANN indexes are partition-based and graph-based indices. Inverted file index (IVF) is a partition-based index and includes many different implementations. ScaNN (Guo et al., 2020) and SOAR (Sun et al., 2024) identify key mathematical properties of ANNS and use them to improve index efficiency. SPANN (Chen et al., 2021) uses disk-based storage to search over billion-scale datasets and supports updates (Xu et al., 2023). FAISS-IVF (Douze et al., 2024) provides an in-memory implementation IVF which can utilize GPUs (Johnson et al., 2019). In this work we show how IVF indexes can be extended with cracking, and our approach is directly applicable to the above. We focus on IVF indexes as they offer the best trade-off between index build time, search performance, and memory overhead, which is critical to minimize overhead in large scale deployments on embedding data lakes, as covered in Section 3. The cracking paradigm can be generalized to other ANN indexes as well. With AV-tree(Lampropoulos et al., 2023) the authors show how to extend a tree-based index with cracking to supports k-NN queries targeting ENN search. Hash-based methods (Andoni et al., 2015; Indyk and Motwani, 1998; Datar et al., 2004; Jain et al., 2008; Wang et al., 2017) and tree-based methods (Muja and Lowe, 2014; Liu et al., 2004; Wang et al., 2013) also broadly fall in the partition-based category. Each index has a unique construction method and data structures. Adopting cracking ideas for them would require solving unique challenges. Finally, graph-based methods (Malkov and Yashunin, 2018; Malkov et al., 2014; Wang et al., 2012; Dong et al., 2011; Wang et al., 2013; Hajebi et al., 2011) currently deliver the state-of-the-art performance in terms of the Queries Per Second vs. Recall trade-off, but this comes at high memory and index construction cost. Thus, graph-based indexes are ill-suited when the goal is minimizing the index construction cost and memory overhead. Cracking is applicable here and could push search performance even further. For example, future research can show how to dynamically change the connections in the graphs based on the arriving query distribution.
@@
Adaptive indexes for specialized data types: Adaptive indexing techniques, such as Dumpy (Wang et al., 2023, 2024) and ADS (Zoumpatianos et al., 2014) are effective in data series similarity search. They are specifically designed for time-series data and include methods like Symbolic Aggregate Approximation (SAX) (Lin et al., 2003) and indexable SAX (iSAX) (Shieh and Keogh, 2008). We focus on ANN indexing methods that are the standard used by RAG systems, and due to their broad applicability across various data modalities which lack sequential relationships. Specialized, adaptive and incrementally constructed indexes are relevant candidates for embedding lake deployments, particularly for the subset of the datasets with inherent sequential dependencies such as energy, weather, financial data, audio, and video (Raza et al., 2015; Mirylenka et al., 2016; Shasha, 1999; Kashino et al., 1999).
@@
ANN index maintenance: Work on index maintenance focuses on mitigating performance degradation caused by data distribution shifts due to data updates (insertions and deletions). DeDrift (Baranchuk et al., 2023) handles data or query distribution shifts keeping the total number of partitions fixed. LIRE (Xu et al., 2023) performs localized split and merge operations when partitions change after insertions/deletions. Ada-IVF (Mohoney et al., 2024) monitors partition size changes from updates and the partition access distribution from the queries. It combines this information to prioritize the maintenance operations. Finally, Quake (Mohoney et al., 2025) is an adaptive index for dynamic workloads with changing access patterns and heavy updates. They use a cost model to decide on maintenance operations, and dynamically set the nprobe parameter. Similar to us, these approaches demonstrate the effectiveness of localized operations for IVF indexes and the advantages of adapting to query distribution. However, they cannot be applied directly for our setting of incremental index construction to minimize cumulative time as the number of queries increases. These works trigger maintenance operations from incoming updates, or focus on distribution shits. Our focus is on the static setting, and we follow the database cracking paradigm: “index maintenance should be a byproduct of query processing, not of updates” (Idreos et al., 2007). Thus, CrackIVF operations are triggered solely from the queries.
7. Conclusion
@@
We introduced CrackIVF, a cracking-based IVF index for Approximate Nearest Neighbor Search (ANNS), aimed to be used with RAG systems over embedding data lakes. By dynamically adapting to increasing query workloads, it eliminates the need for costly up-front indexing, allowing immediate query processing. Our experimental results show that CrackIVF can process over 1 million queries before conventional methods even finish building an index, achieving 10-1000x faster initialization times. This makes it particularly effective for cold data, infrequently accessed datasets, and rapid access to unseen data.
Acknowledgements.